OpenStudy (anonymous):

HELP, it's 4:45 am my time and I have to get up at 6 am for work, my brain is fried.......... Simplify the rational expression 6-2x/4x-12

OpenStudy (lgbasallote):

factor out the numerator and denominator...you'll be able to cancel something

OpenStudy (anonymous):

So far, I've come up with 2(-x+3) / 4(x-3)

OpenStudy (anonymous):

But not sure what to do now

OpenStudy (lgbasallote):

that's good! one trick you should know... (-x+3) = -(x-3) if you pull out a negative sign..you can invert the operations inside a grouping symbol

OpenStudy (anonymous):

so the two would cancel out leaving 2 / 4 ?

OpenStudy (anonymous):

or 1/2

OpenStudy (anonymous):

is that right?

OpenStudy (lgbasallote):

yes! :DDD

OpenStudy (anonymous):

ty sooooooooooooooooooo much

OpenStudy (lgbasallote):

oh wait...you pulled out a negative remember?

OpenStudy (callisto):

Hmm.. it should be a negative number instead of a positive one :(

OpenStudy (callisto):

6-2x/4x-12 = 2(3-x) / 4(x-3) = -2(x-3) / 4(x-3) = -2/4 = -1/2

OpenStudy (lgbasallote):

well i did tell him..maybe he chaged his answer?

Directrix (directrix):